OpenStudy (anonymous):

I do not understand the greatest common factor . This is what the first question it for 24

OpenStudy (amistre64):

you dont understand what a greatest comon factor is? or what it is for a certain problem??

OpenStudy (amistre64):

i suppose one would suggest the other :)

OpenStudy (anonymous):

The greatest common factor is the largest positive integer that divides two numbers without a remainder. It is commonly found by either prime factorisation or the Euclidean algorithm, by uncultured cretins and cultured mathematicians, respectively.

OpenStudy (amistre64):

that algorithm ..is that the one subtracted from another and on down the line?

OpenStudy (anonymous):

Of course.

OpenStudy (anonymous):

yes, ok. so i have to be able to divid then number into correct

OpenStudy (anonymous):

Yes. What numbers are you trying to find it for? IF they are > 10000, look up Euclidean algorithm, if not, factorise them.

OpenStudy (anonymous):

so then like 24. 12,8,6,4,2 are the factor

OpenStudy (anonymous):

Well, they are, but you want the prime factorisation: 24 = 2 * 12 = 2 * 2 * 6 = 2 * 2 * 2 * 3 = 2^3 x 3 It is easier (quicker) to compare numbers this way, though if you found ALL the factors for both, then the highest of them would be the answer, but it's easier to find the prime factorisation and the highest common factor is the part of that which is in both.

OpenStudy (anonymous):

my home work is telling me find the gratest common fator and it gives me a example 36: and the answer is writen down 36,18,9, 12,6,4,3,2,1 and 12 this number is underline so that is telling me that is the greates common factor.

OpenStudy (anonymous):

How old are you?

OpenStudy (anonymous):

If you're young, then you're probably just meant to find all factors and use that. Oh, wait, anyone who does HCF problems is young.

OpenStudy (anonymous):

Quicker to go: 24 = 2^3 x 3 36 = 2^2 x 3^2 HCF = 2^2 x 3 = 12, though

OpenStudy (anonymous):

I am in 5th . does that mean you can not help me?

OpenStudy (anonymous):

so then the answer I gave you 36,18,9,12,6,4,3,2,1, is not correct for 36?

OpenStudy (anonymous):

They are all the factors of 36, yes. 12 is the highest factor of both 24 and 36 so is the answer. You can work out the highest common factor of a number that way I just find it tedious)

OpenStudy (anonymous):

so sorry. but, that was the way my home work is written. like my next problem is 36 and 9 so my answer will be 18,12,9,6,4,2,1?

OpenStudy (anonymous):

I forgot 3 .

OpenStudy (anonymous):

so my greatest common factor would be 3?

OpenStudy (anonymous):

No, because 9 is a factor of 9 (and 36)

OpenStudy (anonymous):

oh, I see I did not think I could use 9 I think I understand. let's try this again. 21 and 28 the common would be 7?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Thank you so much for all your help.